A patient has Lactate Ringer's (LR) solution infusing at 100 mL/hour. The patient consumes 120 mL of juice for breakfast, 240 mL of coffee and 120 mL of ice cream for lunch and 240 mL of milk for dinner. The patient also drank 450 mL of water. Calculate the intake of this patient from 6 am to 6 pm.
Answer #1

intake of the patient`s total hours are = 6 AM to 6 PM = 12 hours

Infused LR solution = 100 ml/hour

Total hour = 12

so total infused LR solution = 12 × 100 = 1200 ml

Patient consumes total liquid =

  • 120 ml juice for break fast
  • 240 ml of coffee and 120 ml of ice cream for lunch
  • 240 ml of milk for dinner
  • 450 ml of water

Total intake = 120+240+120+240+450+1200 = 2370